# Linzhan Mou5, Linqi Song4, Zhenan Sun [^2] Corresponding authors. 1,3, Ying Wei22footnotemark: 2 6 1

### v2

@@ 5 Conclusion
  • Section G: Limitations and broader impacts.
  • Section H: More visualization examples.
@@ A.1 Network Quantization
Network quantization [19, 23, 22, 51] is a widely utilized technique in neural networks aimed at reducing model size and memory usage. Research in this area generally falls into two main categories: quantization-aware training (QAT) [59, 3, 44] and post-training quantization (PTQ) [38, 34, 4, 29]. QAT involves training quantized model weights using additional data, often with the assistance of a straight-through estimator (STE) [5]. However, the computational cost associated with QAT poses challenges, particularly for large language models (LLMs) with millions of parameters, which necessitate significant amounts of data for retraining [33, 14]. In contrast, PTQ has gained popularity for LLMs [48, 21, 58, 55, 35] due to its efficient approach, involving the training of quantized models using a small amount of data, known as calibration data [17]. However, PTQ often leads to significant performance degradation, especially when employing low-bit settings  [17, 47, 41, 28]. Consequently, our work focuses on enhancing the performance of low-bit PTQ quantized models.
@@ A.2 Post Training Quantization of LLM
Post-training quantization for LLMs can be categorized into weight-only quantization [30, 13, 26, 27] and weight-activation quantization [54, 52, 50]. We focus on 4-bit weight-activation quantization due to the actual speedup it provides with low-bit quantization kernels [1]. Quantizing LLMs faces challenges due to activation outlier features persisting across different tokens and layers [12, 48]. Some approaches [12, 61] retain a small portion of crucial outlier channels at high precision (e.g., INT8), which poses challenges to hardware compatibility and leads to additional memory footprint. Other methods [52, 49, 42] attempt to shift quantization difficulty from activation to weight channels. However, the learnable equivalent transformation in OmniQuant [42] and the affine transform matrix in AffineQuant [36] exhibit instability as discussed in Section 2. The channel disassembly and assembly in QLLM [31], coupled with LoRA-tuning, incur significant time costs. Notably, these methods demonstrate poor performance under W4A4 quantization. We attribute this degradation to the ineffective handling of outlier features, especially massive outliers. Hence, we propose DuQuant to effectively eliminate outlier features through rotation matrices and channel permutation, achieving state-of-the-art performance. In contrast with QuaRot [2] also utilizing hadamard matrices to enhance weight-activation quantization, our approach uniquely incorporates knowledge about the actual outlier channels. Furthermore, unlike QuaRot, which relies on GPTQ [17] for weight quantization, our channel permutation has been proven helpful and efficient, facilitating a faster quantization process. The more detailed analysis and comparison with QuaRot are left in Appendix F. In addition, unlike RPTQ [55] and SKVQ [15], which use channel reordering to cluster similar activations, our method employs Permutation transformations with a fundamentally different goal: to evenly distribute outliers across blocks. This balanced distribution is crucial for enabling effective secondary rotations, ultimately leading to smoother activations that facilitate easier quantization.
@@ Appendix C Additional Implementation Details
In this work, all experiments are done on NVIDIA RTX 3090 GPUs for small-scale models and NVIDIA A100 GPUs for large-scale models.
@@ Appendix D More Empirical Results
As LLaMA3 models have proven to be sensitive to quantization, we apply our DuQuant to the LLaMA3-70B and present the results in Table D3. Due to time constraints, we do not add learnable weight clipping. The results demonstrate that our RAP-quantized models outperform SmoothQuant by 12.9% on Commonsense QA tasks and significantly reduce perplexity across the WikiText2, C4, and PTB datasets. These improvements underscore the robustness of our DuQuant method when applied to the LLaMA3-70B model.
@@
To thoroughly evaluate the effectiveness of our DuQuant models, we conduct comprehensive assessments under the W6A6 quantization setting. The perplexity results for language generation tasks are displayed in Table D4, while the zero-shot accuracy for Commonsense QA tasks is detailed in Tables D5 and D6. Our findings reveal that DuQuant not only surpasses other baselines but also achieves nearly lossless performance with FP16 models in these tasks. Interestingly, in several instances, DuQuant slightly outperforms DuQuant+LWC . This suggests that the Rotation and Permutation transformations alone are sufficient to create highly competitive quantized models under W6A6 settings, without the need for additional enhancements such as the learnable weight clipping (LWC) technique. These outcomes highlight the exceptional versatility and robustness of DuQuant across various quantization scenarios, confirming its potential as a leading solution in post-training quantization for large language models.
@@ Appendix E More Ablation Studies
E.1 End-to-end Time Speedup and Memory Saving
In Section 4.2, we analyzed the layer-wise time speedup for the LLaMA2-7B and LLaMA2-13B models. Here, we present comprehensive end-to-end results for time speedup and memory savings achieved with the LLaMA2-7B model on a single NVIDIA RTX 3090 GPU. As shown in Table E7, DuQuant achieves a maximum speedup of 2.01×2.01$\times$2.01 × during the prefilling phase, with speedup increasing as the batch size grows. Additionally, from Table E8, DuQuant demonstrates significant memory savings, effectively reducing memory usage by up to 3.20×3.20$\times$3.20 × through quantization. These results underscore the efficiency of DuQuant in optimizing resource utilization, highlighting its potential to enhance performance and reduce costs in deploying large language models, particularly in resource-constrained environments.
Table E7: End-to-end prefilling speedup on LLaMA2-7B model.
@@ E.1 End-to-end Time Speedup and Memory Saving
Table E8: Peak memory usage during prefilling phase of LLaMA2-7B model.
@@ E.4 Effects of Calibration Datasets
To further explore the robustness of DuQuant under varying calibration conditions, we generate random calibration data within the vocabulary range of the model, setting the sample count to 256. The results, shown in Table E13, indicate that even in calibration-free settings, our method continues to perform well, achieving results that are competitively close to those obtained with actual calibration data. This demonstrates that DuQuant could provide a viable solution in real-world scenarios where obtaining specific calibration data is challenging or impossible. The ability to maintain high performance without traditional calibration data opens avenues for deploying quantized models in environments with strict privacy or data availability limitations, suggesting a promising direction for future research to enhance model adaptability and deployment flexibility.
@@ Appendix F Detailed Comparison with QuaRot
Table F15: Evaluation results between QuaRot and DuQuant under W4A4 quantization.
@@
Experimental results underscore the superiority of DuQuant over QuaRot. For a fair comparison, we reproduce the QuaRot under 4-bit per-channel weight and per-token activation asymmetric quantization. Table F15 displays the perplexity (PPL) and zero-shot accuracy for models LLaMA1-7B, LLaMA2-7B, and LLaMA3-8B. Our DuQuant method consistently outperforms QuaRot-RTN across all benchmarks, showcasing our advanced weight matrix management. Furthermore, compared to QuaRot-GPTQ, DuQuant and DuQuant+LWC achieve better average accuracy across six QA tasks and demonstrate superior performance on the WikiText and C4 datasets, particularly for LLaMA3-8B.
Additionally, we assess the effectiveness of the rotation matrices utilized in DuQuant  which incorporate prior knowledge against the Hadamard matrices used in QuaRot. For a fair comparison, we omit the permutation step in DuQuant and directly contrast it with QuaRot-RTN. Results in Table F16 show that our DuQuant without permutation outperforms QuaRot by a clear margin, which confirms that our rotation transformation is more effective than Hadamard by leveraging prior knowledge. It is worth noting that because a Hadamard matrix is orthogonal and symmetric, it multiplies by itself to yield the identity matrix. In other words, the Hadamard matrix is not suitable for greedy searches aimed at finding smaller outliers. These findings differentiate DuQuant from QuaRot and highlight the effectiveness of our approach in managing outliers for post-training quantization of large language models.
Appendix G Limitations and Broader Impacts

### v3

@@ 5 Conclusion
Acknowledgements
This work is supported in part by the National Natural Science Foundation of China (Grant No. U23B2054, 62276263, and 62371411), the Research Grants Council of the Hong Kong SAR under Grant GRF 11217823 and Collaborative Research Fund C1042-23GF, InnoHK initiative, the Government of the HKSAR, Laboratory for AI-Powered Financial Technologies.
@@ Acknowledgements
  • Section G: Algorithm for rotation matrix.
  • Section H: Limitations and broader impacts.
  • Section I: More visualization examples.
@@ A.1 Network Quantization
Neural networks achieve great success in recent years [32, 66, 72, 39, 62, 19, 63], quantization [1, 64, 68, 76, 58, 77] is a widely utilized technique aimed at reducing model size and memory usage. Research in this area generally falls into two main categories: quantization-aware training (QAT) [73, 3, 50] and post-training quantization (PTQ) [42, 61, 75]. QAT involves training quantized model weights using additional data, often with the assistance of a straight-through estimator (STE) [5]. However, the computational cost associated with QAT poses challenges, particularly for large language models (LLMs) with millions of parameters, which necessitate significant amounts of data for retraining [37, 16, 9]. In contrast, PTQ has gained popularity for LLMs [55, 23, 70, 59, 38, 21] due to its efficient approach, involving the training of quantized models using a small amount of data, known as calibration data [18]. However, PTQ often leads to significant performance degradation, especially when employing low-bit settings  [18, 54, 47, 31]. Consequently, our work focuses on enhancing the performance of low-bit PTQ quantized models.
@@ A.2 Post Training Quantization of LLM
Post-training quantization for LLMs can be categorized into weight-only quantization [33, 14, 29, 30] and weight-activation quantization [69, 65, 76]. We focus on 4-bit weight-activation quantization due to the actual speedup it provides with low-bit quantization kernels [1]. Quantizing LLMs faces challenges due to activation outlier features persisting across different tokens and layers [13, 65]. Some approaches [13, 76] retain a small portion of crucial outlier channels at high precision (e.g., INT8), which poses challenges to hardware compatibility and leads to additional memory footprint. Other methods [65, 60, 48] attempt to shift quantization difficulty from activation to weight channels. However, the learnable equivalent transformation in OmniQuant [48] and the affine transform matrix in AffineQuant [40] exhibit instability as discussed in Section 2. The channel disassembly and assembly in QLLM [35], coupled with LoRA-tuning, incur significant time costs. Notably, these methods demonstrate poor performance under W4A4 quantization. We attribute this degradation to the ineffective handling of outlier features, especially massive outliers. Hence, we propose DuQuant to effectively eliminate outlier features through rotation matrices and channel permutation, achieving state-of-the-art performance. In contrast with QuaRot [2] also utilizing Hadamard matrices to enhance weight-activation quantization, our approach uniquely incorporates knowledge about the actual outlier channels. Furthermore, unlike QuaRot, which relies on GPTQ [18] for weight quantization, our permutation transformation has been proven helpful and efficient, facilitating a faster quantization process. The more detailed analysis and comparison with QuaRot are left in Appendix F. In addition, unlike RPTQ [70] and SKVQ [17], which use channel reordering to cluster similar activations, our method employs Permutation transformations with a fundamentally different goal: to evenly distribute outliers across blocks. This balanced distribution is crucial for enabling effective secondary rotations, ultimately leading to smoother activations that facilitate easier quantization.
@@ Appendix C Additional Implementation Details
In this work, all experiments are done on NVIDIA RTX 3090 GPUs for small-scale models and NVIDIA A100 GPUs for large-scale models. We set sequence length to 2048 for all evaluation tasks.
@@ Appendix D More Empirical Results
Long-context Evaluation Results.
LongBench [4] is proposed to access the long-context generation ability of LLMs, which covers several key long-text application scenarios. We evaluate the 4-bit quantized Vicuna models on five different tasks. Specifically, Qasper, MultiFieldQA, and NarrativeQA (F1 score) are Single-Document QA tasks; DuReader (Rouge-L score) and 2WikiMultihopQA (F1 score) are Multi-Document QA tasks; QMSum, GovReport (F1 score) and MultiNews (Rouge-L score) are Summarization tasks; TREC (Accuracy CLS), TriviaQA (F1 score), and SAMSum (Rouge-L score) are Few-shot Learning tasks; and RepoBench-P (similarity score) is Code Completion task. Table D3 and Table D4 show that our DuQuant outperforms other baselines by a clear margin, maintaining the ability for long context generation tasks compared with FP16 models.
Table D3: Long-context generation results on the LongBench benchmark for 4-bit Vicuna-v1.5-7B.
Vicuna-v1.5-7B RepoBench-P MultiFieldQA-en GovReport MultiNews DuReader 2WikiMQA TriviaQA
FP16 48.23 38.30 27.93 26.91 25.53 18.02 82.59
SmoothQuant 25.92 4.66 2.62 6.05 4.24 2.02 1.62
OmniQuant 14.97 2.30 2.51 2.64 1.87 0.48 0.81
Atom 29.34 31.15 23.60 24.60 19.41 17.10 67.20
DuQuant 47.66 35.62 25.66 25.85 23.15 15.09 78.91
Vicuna-v1.5-7B QMSum MultiFieldQA-zh NarrativeQA Qasper SAMSum TREC Avg
FP16 21.07 32.56 14.96 23.27 41.06 66.00 35.88
SmoothQuant 2.00 0.88 1.75 4.11 1.55 15.00 5.57
OmniQuant 3.93 1.40 1.10 1.62 0.61 1.00 2.71
Atom 20.24 21.55 11.57 17.97 37.94 58.00 29.21
DuQuant 21.15 29.56 11.31 19.98 42.24 64.00 33.86
Table D4: Long-context generation results on the LongBench benchmark for 4-bit Vicuna-v1.5-13B.
Vicuna-v1.5-13B RepoBench-P MultiFieldQA-en GovReport MultiNews DuReader 2WikiMQA TriviaQA
FP16 43.08 42.69 28.43 26.53 27.57 29.40 86.81
SmoothQuant 11.57 1.64 2.81 3.54 6.71 1.39 1.83
OmniQuant 8.46 4.32 0.74 2.83 13.83 0.75 1.13
Atom 37.31 37.31 19.34 23.39 21.79 15.16 80.75
DuQuant 38.09 44.12 26.97 26.59 26.02 22.07 83.04
Vicuna-v1.5-13B QMSum MultiFieldQA-zh NarrativeQA Qasper SAMSum TREC Avg
FP16 21.24 40.44 15.41 24.41 41.97 68.00 40.64
SmoothQuant 2.95 0.82 0.97 2.18 0.35 1.50 4.21
OmniQuant 1.78 1.06 0.62 0.68 0.45 9.00 4.58
Atom 20.23 28.02 8.81 17.67 38.72 59.00 33.58
DuQuant 20.72 30.85 13.36 18.93 42.67 66.50 38.13
Table D5: More results on MT-Bench.
DuQuant v.s. FP16 Former Win Tie Former Loss Vicuna-v1.5-7B 36 56 68 Vicuna-v1.5-13B 43 53 64
Comparison with FP16 models on MT-Bench.
We conducted additional comparisons using the MT-Bench between our INT4 quantized models and the FP16 models. As shown in Table D5, for both 7B and 13B models, our DuQuant performs comparably to FP16, which further underscores the effectiveness of dual transformations in maintaining high accuracy even with reduced precision.
Results for 4-bit Mistral-7B and Phi2-2.8B.
We have extended the application of DuQuant to include Mistral [28] and Phi2 [27] under 4-bit WA quantization. From Table D6, we can observe that DuQuant largely surpasses other baselines, particularly with Mistral-7B. Regarding the Phi2-2.8B model, it often experiences instability in matrix multiplication between queries and values, leading to overflow issues and posing great challenges to quantization. However, while DuQuant may not perform as well as FP models, it still significantly outperforms other baselines. In addition, we have visualized the massive outliers in the down projection layer of the Mistral-7B model and the feature space after our dual transformations. These visualizations are shown in Figure I9. It can be observed that our DuQuant perfectly eliminates these outliers. These results underscore the effectiveness of our dual transformation approach in addressing massive outliers across various types of LLMs.
Table D6: Perplexity results of Mistral-7B and Phi2-2.8B under 4-bit weight-activation quantization.
Model Method WikiText2 C4 Model Method WikiText2 C4
Mistral-7B W4A4 FP16 5.25 7.75 Phi2-2.8B W4A4 FP16 9.71 12.76
RTN 306.26 300.07 RTN 230.59 253.79
SmoothQuant 100.59 158.02 SmoothQuant 63.84 83.24
OmniQuant 5490.31 6094.82 OmniQuant NaN NaN
Atom 8.65 12.43 Atom 35.72 41.26
DuQuant 5.86 8.48 DuQuant 20.65 22.49
@@
As LLaMA3 models have proven to be sensitive to quantization, we apply our DuQuant to the LLaMA3-70B and present the results in Table D7. Due to time constraints, we do not add learnable weight clipping. The results demonstrate that our DuQuant-quantized models outperform SmoothQuant by 12.9% on Commonsense QA tasks and significantly reduce perplexity across the WikiText2, C4, and PTB datasets. These improvements underscore the robustness of our DuQuant method when applied to the LLaMA3-70B model.
@@
To thoroughly evaluate the effectiveness of our DuQuant models, we conduct comprehensive assessments under the W6A6 quantization setting. The perplexity results for language generation tasks are displayed in Table D8, while the zero-shot accuracy for Commonsense QA tasks is detailed in Tables D9 and D10. Our findings reveal that DuQuant not only surpasses other baselines but also achieves nearly lossless performance with FP16 models in these tasks. Interestingly, in several instances, DuQuant slightly outperforms DuQuant+LWC . This suggests that the Rotation and Permutation transformations alone are sufficient to create highly competitive quantized models under W6A6 settings, without the need for additional enhancements such as the learnable weight clipping (LWC) technique. These outcomes highlight the exceptional versatility and robustness of DuQuant across various quantization scenarios, confirming its potential as a leading solution in post-training quantization for large language models.
@@ Appendix E More Ablation Studies
E.1 Time Speedup and Memory Saving
The current generation of LLMs usually splits into pre-filling and decoding phases and deploys on two separate machines [44]. Here, we present more speedup and memory-saving results for these two phases achieved with the LLaMA2-7B model on a single NVIDIA RTX 3090 GPU. We set the input sequence length to 2048 and the decoding steps to 256. End-to-end results for time speedup and memory savings during the pre-filling stage are shown in Table E11 and E12. We can observe that DuQuant achieves a maximum speedup of 2.01×2.01$\times$2.01 × during the pre-filling phase, with speedup increasing as the batch size grows. From Table E12, DuQuant demonstrates significant memory savings, effectively reducing memory usage by up to 3.20×3.20$\times$3.20 × through quantization. For the decoding phase, we enlarge the batch size to 64 and measure speedup along with memory usage for one LLaMA2-7B layer, constrained by the 24 GB memory of the GPU. As shown in Table E13, DuQuant maintains speedup and memory usage comparable to QuaRot. These results underscore the efficiency of DuQuant in optimizing resource utilization, highlighting its potential to enhance performance and reduce costs in deploying large language models, particularly in resource-constrained environments.
Table E11: End-to-end pre-filling speedup on LLaMA2-7B model.
@@ E.1 Time Speedup and Memory Saving
Table E12: Peak memory usage during pre-filling phase of LLaMA2-7B model.
@@
Table E13: Decoding phase results of one LLaMA2-7B layer with a batch size of 64.
Method Time (ms) Saving Factor Memory (GB) Saving Factor
FP16 659 - 3.550x -
SmoothQuant 437 1.508x 1.669 2.127x
QLLM OOM - OOM -
QuaRot 457 1.442x 1.678 2.116x
DuQuant 499 1.321x 1.677 2.117x
@@ E.4 Effects of Calibration Datasets
To further explore the robustness of DuQuant under varying calibration conditions, we generate random calibration data within the vocabulary range of the model, setting the sample count to 256. The results, shown in Table E18, indicate that even in calibration-free settings, our method continues to perform well, achieving results that are competitively close to those obtained with actual calibration data. This demonstrates that DuQuant could provide a viable solution in real-world scenarios where obtaining specific calibration data is challenging or impossible. In addition, our findings suggest that outliers are inherent to certain model layers, reflecting characteristics of the model weights or modules, especially in recent LLMs. This aligns with concurrent research: [67] identified consistent massive outliers at the FFN down projection layer in GLU-based LLMs, such as LLaMA, Mistral, Mixtral, SOLAR, and Gemma, while [43] reported that, although OPT models are sensitive to different calibration sets, newer models demonstrate robustness to outliers and maintain stable activations. These insights reinforce the idea that outliers are more linked to the internal structure of model weights and modules than to calibration data. DuQuant’s ability to deliver high performance without relying on traditional calibration data creates opportunities for deploying quantized models in environments with stringent privacy requirements or limited data availability. This highlights a promising direction for future research, focusing on improving model adaptability and deployment flexibility.
@@ Appendix F Detailed Comparison with QuaRot
Table F20: Comparison of quantization settings between QuaRot and DuQuant.
Setting Weight Activation Query
QuaRot per-channel symmetric per-token symmetric FP16
DuQuant per-channel asymmetric per-token asymmetric per-token asymmetric
Table F21: Evaluation results between QuaRot and DuQuant under DuQuant settings.
@@
Table F22: Evaluation results between QuaRot and DuQuant under QuaRot settings.
Model Method WikiText2 $\downarrow$ C4 $\downarrow$ PIQA WinoGrande HellaSwag ARC-E ARC-C LAMBADA Avg $\uparrow$
LLaMA2-7B W4A4 QuaRot Setting FP16 5.47 6.97 79.11 69.06 75.99 74.58 46.25 73.90 69.82
QuaRot-RTN 8.37 - 72.09 60.69 65.4 58.88 35.24 57.27 58.26
QuaRot-GPTQ 6.1 - 76.77 63.77 72.16 69.87 40.87 70.39 65.64
DuQuant 6.23 7.91 76.28 66.93 72.96 69.99 40.53 69.61 66.05
DuQuant+LWC 6.01 7.67 77.64 67.8 72.97 70.37 41.81 69.53 66.69
LLaMA2-13B W4A4 QuaRot Setting FP16 4.88 6.46 80.47 72.22 79.39 77.48 49.23 76.75 72.59
QuaRot-RTN 6.09 - 77.37 67.32 73.11 70.83 43.69 70.66 67.16
QuaRot-GPTQ 5.4 - 78.89 70.24 76.37 72.98 46.59 73.67 69.79
DuQuant 5.39 7.05 78.51 70.88 76.80 74.62 48.21 73.92 70.49
DuQuant+LWC 5.27 6.93 78.73 70.88 77.20 74.07 47.27 73.96 70.35
@@
Experimental results underscore the superiority of DuQuant over QuaRot. We first summarize the experimental setting differences between the original paper of QuaRot with ours in Table F20***Following prior works [48, 35, 40], we describe operations on Query, Key, and Value states as per-tensor quantization for consistency. To be precise, these operations are effectively applied on a per-head basis.. For a fair comparison, we reproduce the QuaRot under 4-bit per-channel weight and per-token activation asymmetric quantization. Table F21 displays the perplexity (PPL) and zero-shot accuracy for models LLaMA1-7B, LLaMA2-7B, and LLaMA3-8B. Our DuQuant method consistently outperforms QuaRot-RTN across all benchmarks, showcasing our advanced weight matrix management. Furthermore, compared to QuaRot-GPTQ, DuQuant and DuQuant+LWC achieve better average accuracy across six QA tasks and demonstrate superior performance on the WikiText and C4 datasets, particularly for LLaMA3-8B. Moreover, we further provide the results of DuQuant under the setting utilized in the original paper of QuaRot in Table F22. DuQuant still surpasses QuaRot by a large margin.
Additionally, we assess the effectiveness of the rotation matrices utilized in DuQuant  which incorporate prior knowledge against the Hadamard matrices used in QuaRot. We omit the permutation step in DuQuant and directly contrast it with QuaRot-RTN. Results in Table F23 show that our DuQuant without permutation outperforms QuaRot by a clear margin, which confirms that our rotation transformation is more effective than Hadamard by leveraging prior knowledge. It is worth noting that because a Hadamard matrix is orthogonal and symmetric, it multiplies by itself to yield the identity matrix. In other words, the Hadamard matrix is not suitable for greedy searches aimed at finding smaller outliers. These findings differentiate DuQuant from QuaRot and highlight the effectiveness of our approach in managing outliers for post-training quantization of large language models.
Appendix G Algorithm for Rotation Matrix
Algorithm 1 Construction of the Rotation Matrix
Input: Pre-initialized rotation matrix $\tilde{\mathbf{R}}$, greedy search steps $N$, activation matrix $\mathbf{X}$ with shape of [$T,C_{\text{in}}$]
Output: Rotation matrix $\hat{\mathbf{R}}$
function get_rotation_matrix ($\mathbf{X},\tilde{\mathbf{R}},N$)
1:$T$, $C_{\text{in}}$ = $\mathbf{X}.\text{shape}$
2:$\mathbf{R}$ = eye($C_{\text{in}}$)
3:$a=\max_{i,j}|\mathbf{X}_{ij}|$
4:for $k$ in $1,\ ...,N$ do
[…]
7:    Obtain randomly initialized orthogonal matrix $\mathbf{Q}^{\prime}$ with shape of [$C_{\text{in}}-1$, $C_{\text{in}}-1$]
8:    $\mathbf{Q}^{\prime}$ = concat([zeros$(C_{\text{in}}-1,1)$, $\mathbf{Q}^{\prime}$ ], dim=1)
9:    $\mathbf{Q}$ = concat([zeros$(1,C_{\text{in}})$, $\mathbf{Q}^{\prime}$], dim=0)
10:    $\mathbf{Q}[0,0]=1$
11:    $\mathbf{R}^{\prime}$ = matmul($\tilde{\mathbf{R}}$, $\mathbf{Q}$)
12:    $\mathbf{R}^{\prime}$[:, outlier_channel], $\mathbf{R}^{\prime}$[:, 0] = $\mathbf{R}^{\prime}$[:, 0], $\mathbf{R}^{\prime}$[:, outlier_channel]
13:    $\mathbf{R}^{\prime}$[outlier_channel, :], $\mathbf{R}^{\prime}$[0, :] = $\mathbf{R}^{\prime}$[0, :], $\mathbf{R}^{\prime}$[outlier_channel]
14:    $\mathbf{R}$ = matmul($\mathbf{R}$, $\mathbf{R}^{\prime}$)
15:    ${\mathbf{X}}$ = matmul(${\mathbf{X}}$, $\mathbf{R}^{\prime}$)
16:    if $\max_{i,j}|\mathbf{X}_{ij}|<a$ then
17:         $\hat{\mathbf{R}}=\mathbf{R}$
18:         $a=\max_{i,j}|\mathbf{X}_{ij}|$
19:    end if
20:end for
21:return $\hat{\mathbf{R}}$
Appendix H Limitations and Broader Impacts